Use the point slope form to write an equation of the line that passes through the given point and has the given slope.
(5,-7), m=3/5

Answer by stanbon(75887)

Use the point slope form to write an equation of the line that passes through the given point and has the given slope.
(5,-7), m=3/5
======
y+7 = (3/5)(x-5)
----
y = (3/5)x -3-7
----
y = (3/5)x - 10
=====================
Cheers,
Stan H.
=================